Acme Brick Company estimated its annual total cost function to be: Y = $320,000 + 2.56x...

  1. Acme Brick Company estimated its annual total cost function to be:

Y = $320,000 + 2.56x

Assuming that Y represents total cost and x equals the number of units sold, use this equation to answer the following questions:

What is the firm’s total fixed cost?

What is the firm’s variable cost per unit?

Compute total costs if the firm sells 10,000 units.

Compute total costs if the firm sells 30,000 units.
Answer #1

Total cost function is Y = $320,000+2.56x

Y - Total cost

x - No of units sold

Firms total fixed cost = $320,000

(Fixed cost is constant and will not change by change in the sales)

Firms variable cost per unit = $2.56

( Variable cost depends on sales quantity i.e., total variable cost = 2.56x)

If the firm sells 10,000 units total cost is

​​​Total cost (Y) = 320,000+(10,000*2.56)

= 320,000+25,600

= $345,600

If the firm sells 30,000 units total cost is

Total cost (Y) = 320,000+(30,000*2.56)

= 320,000+76,800

= $396,800